A bottle with a volume of 198 U. S. fluid gallons is filled at the rate of 1.7 g/min. (Water has a density of 1000 kg/m3, and 1 U.S. fluid gallon = 231 in.3.) How long does the filling take?




1
Expert's answer
2021-09-20T10:00:30-0400

Let the volume of the bottle be "V" and the rate of filling "r=1.7g\/min". The mass of this water is then given as follows:


"m = \rho V"


where "\rho=1000kg/m^2 = 1g/cm^3".

Then the time required to fill the bottle is:


"t = \dfrac{m}{r} = \dfrac{\rho V}{r}"

Converting volume to "cm^3", obtain:


"V = 198\times 231in^3 = 45738\space in^3 = 749511.53\space cm^3"

Finally:


"t = \dfrac{1g/sm^3\cdot 749511.53\space cm^3}{1.7g/min} \approx 440889min\approx 306\space days"

Answer. "306\space days".